OpenStudy (tshimp0629):

If the endpoints of have the coordinates A(9, 8) and B(-1, -2), what is the midpoint of ? (5, 3) (4, 5) (5, 5) (4, 3)

OpenStudy (jiteshmeghwal9):



OpenStudy (jiteshmeghwal9):

midpoint formula\[\left( \frac{x_2+x_1}{2},\frac{y_1+y_2}{2} \right)\]